OpenStudy (ammarah):

Factor the polynomial completely. x^3+64

OpenStudy (solomonzelman):



OpenStudy (imtiaz7):

x^3+64 =(x)^3+(4)^3 By using the cube formula: a^+b^3= (a+b)(a^2-ab+b^2) so here we have a=x and b=3 (x+3)(x^2-3x+9)=0 either x+3=0 or (x^2-3x+9)=0 x=-3 or here a=1, b=-3 and c=9 here we use quadratic formula which is given below and just put the values and solve it: